OpenStudy (anonymous):

I don't understand how cos(110) deg= -cos(30)deg. What does -cos even mean?? I was given a formula cos x= -cos(180-x) which I don't fully understand either.

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow cos = {adjacent \over hypotenuse} }\)

OpenStudy (anonymous):

I understand what cosine is but what does the negative sign do to it??

Parth (parthkohli):

@lalaly @Mertsj

OpenStudy (lalaly):

cos (60) for example=1/2 cos (180-60)=cos(120) =-1/2